Glucose breaks down into elements by the unbalanced formula C6H12O6 → C + H + O. Given a sample of 0.280 moles of Glucose, how many moles of Carbon will be collected among the products? (Remember to balance the equation.)

A. 0.280 mol C
B. 1.00 mol C
C. 1.12 mol C
D. 1.68 mol C
E. 6.00 mol C
1
Expert's answer
2018-01-13T10:40:07-0500
C6H12O6 → 6C + 12H + 6O
n(C) = 6* n(C6H12O6) = 6*0.28mol =1.68 mol
D. 1.68 mol C